OpenStudy (anonymous):

find derivative of y = (e^x − e^−x)/6

OpenStudy (anonymous):

$$\sinh x=\frac12(e^x-e^{-x})\implies y=\frac13\sinh x$$

OpenStudy (anonymous):

Do you know the derivative of \(\sinh x\)?

OpenStudy (anonymous):

why are you doing it with sine ?

OpenStudy (anonymous):

@oldrin.bataku I havent seen the sinh , i was thinking using the regular derivative formulas but I just cant seem to get the right answer

OpenStudy (loser66):

@camilasanchez if you don't know about sinh, cosh. that's ok, just take derivative as usual

OpenStudy (loser66):

try it. and show me your work, please

OpenStudy (anonymous):

@camilasanchez pull out the (1/6) -- easy. What's the derivative of \(e^x\)? Surely just \(e^x\)! What about \(e^{-x}\) -- ah, use the chain rule!

OpenStudy (anonymous):

with the chain rule in e^-x be: -1(e^-2x)(-1) ?

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

No it looks like you're trying to do the power rule!

OpenStudy (anonymous):

$$\frac{d}{dx}e^{-x}=\frac{d(e^{-x})}{d(-x)}\cdot\frac{d(-x)}{dx}=e^{-x}\cdot-1=-e^{-x}$$

OpenStudy (anonymous):

perfect thanks ! thats why I got it wrong I didnt know how to apply the chain rule correctly ! thanks!

OpenStudy (anonymous):

no prob!